OpenStudy (anonymous):

what percent of 86 is 50 to the nearest whole percent 43 percent 58 percent 59 percent 172 percent

OpenStudy (anonymous):

\[\frac{ x }{ 100 } \times 86=50\]

OpenStudy (anonymous):

do you understand what i have done??

OpenStudy (anonymous):

yeas what would happen nect

OpenStudy (jdoe0001):

next, is solve by "x" :|

OpenStudy (jdoe0001):

or isolate "x", x=?

OpenStudy (anonymous):

first of all: divide both sides of the equation by 86 after, multiply both sides of the equation by 100 and what do you get??

OpenStudy (anonymous):

explain using numbers

OpenStudy (anonymous):

if you divide both sides by 86, \[\frac{ x }{ 100 } = \frac{ 50 }{ 86 }\]

OpenStudy (anonymous):

so can you carry out the next step??

OpenStudy (anonymous):

No can you just go through the problem.

OpenStudy (anonymous):

In numbers don't give me the answers.

OpenStudy (anonymous):

i am not giving you answers....you are the one who wants me to go through every step!

OpenStudy (anonymous):

if you multiply both sides by 100, \[x=\frac{ 50 }{ 86 } \times 100\] \[x=??\]

OpenStudy (anonymous):

do you get it now??